# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-B Proof of Theorem 1
Since $F(\boldsymbol{x}^{t})\geq F(\boldsymbol{x}^{*})$ always holds, we drop the negative term $-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ in (137) to obtain
$\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq(1+\frac{a_{t}}{m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\Phi_{t}$ (139)
$\textstyle\leq\left(\prod_{t=0}^{T}(1+a_{t})\right)\left(\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=0}^{T}\Phi_{t}\right).$
@@
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}\!-\!\boldsymbol{x}^{*}\|^{2}]\leq e^{\frac{s}{s-1}}\left(\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\sum_{t=0}^{T}\Phi_{t}\right).$ (140)
@@
[…]
@@
$\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq e^{\frac{s}{s-1}}\left(\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\frac{c_{6}}{m}\right).$ (143)
We proceed to sum both sides of (137) from 00 to $T$ ($T$ can be any positive integer):
$\textstyle\sum_{t=0}^{T}\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq-\sum_{t=0}^{T}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\quad+\sum_{t=0}^{T}(1+\frac{a_{t}}{m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=0}^{T}\Phi_{t}.$ (144)
@@
$\textstyle\sum_{t=0}^{T}(1+\frac{a_{t}}{m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]-\sum_{t=0}^{T}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\leq\frac{a_{0}}{m}\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=1}^{T}\frac{a_{t}}{m}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]-\mathbb{E}[\|\boldsymbol{\bar{x}}^{T+1}-\boldsymbol{x}^{*}\|^{2}]\leq\frac{c_{7}}{m},$ (145)
with $\frac{c_{7}}{m}=(\frac{se^{\frac{s}{s-1}}}{m(s-1)}+1+\frac{1}{m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\frac{c_{6}s}{m(s-1)}$. Here, we have used (141) and (143) in the last inequality.
Substituting (142) and (145) into (144) yields the inequality $\sum_{t=0}^{T}2\lambda_{x}^{t}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq c_{6}+c_{7}$. Further using $\lambda_{x}^{T}\leq\lambda_{x}^{t}$ holding for any $t\in[0,T]$, we arrive at
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq\frac{c_{6}+c_{7}}{2\lambda_{x}^{0}(T+1)^{1-v_{x}}},$ (146)
which implies (5).
@@
To prove the convergence of Algorithm 1 for nonconvex objective functions, we first need to estimate an upper bound on $\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$.
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]=\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*}),$
@@
$\textstyle\quad+\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
×$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle].$ (147)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*}),\!\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\!\boldsymbol{1}_{m}\!\otimes\!g(\boldsymbol{x}^{t}))\!-\!\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*})\|^{2}]+\frac{b_{1}}{\kappa(t+1)^{\beta_{1}}}+\frac{L_{F}b_{2}}{(t+1)^{\beta_{2}}},$ (148)
where $b_{1}$ and $b_{2}$ are the same as those given in (61) and $L_{F}$ is from Lemma 2.
We characterize the second term on the right hand side of (147) by estimating an upper bound on the term $\mathbb{E}[\|\Xi_{x}^{\prime t}-\Xi_{x}^{\prime t-1}\|]$ with $\Xi_{x}^{\prime t}\triangleq\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}\nabla F(\boldsymbol{\bar{x}}^{t})$. By using an argument similar to the derivation of (88), we have
@@
$\textstyle\leq\mathbb{E}[\|(\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t}))^{\top}\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|(\nabla\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1}))^{\top}\nabla F(\boldsymbol{\bar{x}}^{t-1})\|^{2}]$
$\textstyle\quad+2\mathbb{E}[\|(\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t}))^{\top}\nabla F(\boldsymbol{\bar{x}}^{t})\|]$
$\textstyle\quad+2\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{\bar{x}}^{t-1})\|^{2}\|\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1})\|^{2}]$
@@
$\textstyle\quad+2L_{g}^{2}L_{F}^{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}\|^{2}].$ (149)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
×$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*})\|^{2}]$
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$ (152)
$\textstyle\leq\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*})\|^{2}]$
@@
$\textstyle\mathbb{E}[F(\boldsymbol{\bar{x}}^{t+1})-F(\boldsymbol{\bar{x}}^{t})]\leq\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{\bar{x}}^{t}\rangle]$ (153)
$\textstyle\quad+\frac{L_{F}^{2}}{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{\bar{x}}^{t}\|^{2}].$
According to the relation $\boldsymbol{\bar{x}}^{t+1}=\boldsymbol{\bar{x}}^{t}+\boldsymbol{\bar{\vartheta}}_{w}^{t}-\frac{\lambda_{x}^{t}}{m}\boldsymbol{u}^{t}$, which holds due to $\Omega=\mathbb{R}^{n}$ given in the theorem statement, the first term on the right hand side of (153) satisfies
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{\bar{x}}^{t}\rangle]=-\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\frac{\lambda_{x}^{t}}{m}\boldsymbol{u}^{t}\rangle]$
$\textstyle\leq\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]-\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\nabla F(\boldsymbol{\bar{x}}^{t})\rangle]$
$\textstyle\quad-\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\nabla F(\boldsymbol{\bar{x}}^{t})\rangle].$ (154)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\nabla F(\boldsymbol{\bar{x}}^{t})\rangle]$ (155)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]+\frac{1}{\kappa}\mathbb{E}[\|\boldsymbol{\breve{u}}^{t}-\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}].$
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\breve{u}}^{t}-\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]$
$\textstyle\leq(2\bar{L}_{f}^{2}(1+L_{g}^{2})(1+2L_{g}^{2})+4\bar{L}_{g}^{2}L_{f}^{2})\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{\bar{x}}^{t}\|^{2}]$
$\textstyle\leq\frac{c_{8}}{(t+1)^{2\varsigma_{x}}},$ (156)
where $c_{8}$ is given by $c_{8}=C_{x}(2\bar{L}_{f}^{2}(1+L_{g}^{2})(1+2L_{g}^{2})+4\bar{L}_{g}^{2}L_{f}^{2})$ with $C_{x}$ given in (50). Here, we have used $\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{\bar{x}}^{t}\|^{2}]\leq\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\leq\frac{C_{x}}{(t+1)^{2\varsigma_{x}}}$ in the last inequality.
@@
$\textstyle\mathbb{E}[\langle\nabla F(\boldsymbol{\bar{x}}^{t}),\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{\bar{x}}^{t}\rangle]\leq-\frac{\lambda_{x}^{t}}{2m}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]$ (157)
$\textstyle\quad+\frac{b_{5}\lambda_{x}^{t}}{m\kappa(t+1)^{\beta_{1}}}+\frac{b_{6}^{\prime}\lambda_{x}^{t}}{m(t+1)^{\beta_{3}}}+\frac{c_{8}}{\kappa(t+1)^{2\varsigma_{x}}}.$
@@
$\mathbb{E}[F(\boldsymbol{\bar{x}}^{t+1})-F(\boldsymbol{\bar{x}}^{t})]\leq-\frac{\lambda_{x}^{t}}{2m}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]+\Phi_{t},$ (158)

### v3

@@ VII-B Proof of Theorem 1
Since $F(\boldsymbol{x}^{t})\geq F(\boldsymbol{x}^{*})$ always holds, we drop the negative term $\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ in (131) to obtain
$\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq(1+\frac{a_{t}}{m})\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\Phi_{t}$ (133)
$\textstyle\leq\left(\prod_{t=0}^{T}(1+a_{t})\right)\left(\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=0}^{T}\Phi_{t}\right).$
@@
$\vskip-1.99997pt\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}\!-\!\boldsymbol{x}^{*}\|^{2}]\leq e^{\frac{s}{s-1}}\left(\mathbb{E}[\|\mathbf{\bar{x}}^{0}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\sum_{t=0}^{T}\Phi_{t}\right).$ (134)
@@
$\vskip-1.99997pt\textstyle\sum_{t=0}^{T}\frac{1}{(t+1)^{s}}\leq 1+\int_{1}^{\infty}\frac{1}{x^{s}}dx\leq\frac{s}{s-1},$ (135)
@@
$\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq e^{\frac{s}{s-1}}\left(\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\frac{c_{6}}{m}\right).$ (137)
We proceed to sum both sides of (131) from 0 to $T$ ($T$ can be any positive integer):
$\textstyle\sum_{t=0}^{T}\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq-\sum_{t=0}^{T}\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\quad+\sum_{t=0}^{T}(1+\frac{a_{t}}{m})\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=0}^{T}\Phi_{t}.$ (138)
@@
$\textstyle\sum_{t=0}^{T}(1+\frac{a_{t}}{m})\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]-\sum_{t=0}^{T}\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\leq\frac{a_{0}}{m}\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\sum_{t=1}^{T}\frac{a_{t}}{m}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]-\mathbb{E}[\|\mathbf{\bar{x}}^{T+1}-\boldsymbol{x}^{*}\|^{2}]\leq\frac{c_{7}}{m},$ (139)
with $\frac{c_{7}}{m}=(\frac{se^{\frac{s}{s-1}}}{m(s-1)}+1+\frac{1}{m})\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}]+\frac{c_{6}s}{m(s-1)}$. Here, we have used (135) and (137) in the last inequality.
Substituting (136) and (139) into (138) yields the inequality $\sum_{t=0}^{T}2\lambda_{x}^{t}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq c_{6}+c_{7}$.
By using (44) and Assumption 1, we have
$\textstyle\sum_{t=0}^{T}2\lambda_{x}^{t}\mathbb{E}[F(\mathbf{x}_{i}^{t})-F(\boldsymbol{x}^{*})]=\textstyle\sum_{t=0}^{T}2\lambda_{x}^{t}\mathbb{E}\left[F(\mathbf{x}_{i}^{t})-F(\mathbf{\bar{x}}^{t}),\right$.
$\left.\textstyle\quad+F(\mathbf{\bar{x}}^{t})-F(\boldsymbol{x}^{t})\right]+\textstyle\sum_{t=0}^{T}2\lambda_{x}^{t}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$
$\textstyle\leq\sum_{t=0}^{T}4\lambda_{x}^{t}L_{f}\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]+c_{6}+c_{7}$
$\textstyle\leq\frac{4\lambda_{0}^{t}L_{f}\sqrt{\bar{C}_{x}}(v_{x}+\varsigma_{x})}{v_{x}+\varsigma_{x}-1}+c_{6}+c_{7}\triangleq c_{8}.$
Further using $\lambda_{x}^{T}\leq\lambda_{x}^{t}$ holding for any $t\in[0,T]$, we have $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[F(\mathbf{x}_{i}^{t})-F(\boldsymbol{x}^{*})]\leq\frac{c_{8}}{2\lambda_{x}^{0}(T+1)^{1-v_{x}}}$.
@@
To prove the convergence of Algorithm 1 for nonconvex objective functions, we first need to estimate an upper bound on $\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$.
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]=\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*}),$
@@
$\textstyle\quad+\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle].$ (140)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*}),\!\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\!\boldsymbol{1}_{m}\!\otimes\!g(\boldsymbol{x}^{t}))\!-\!\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})\!-\!\nabla F(\boldsymbol{x}^{*})\|^{2}]+\frac{b_{1}}{\kappa(t+1)^{\beta_{1}}}+\frac{L_{F}b_{2}}{(t+1)^{\beta_{2}}},$ (141)
where $b_{1}$ and $b_{2}$ are the same as those given in (55).
We characterize the second term on the right hand side of (140) by estimating an upper bound on the term $\mathbb{E}[\|\Xi_{x}^{\prime t}-\Xi_{x}^{\prime t-1}\|]$ with $\Xi_{x}^{\prime t}\triangleq\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}\nabla F(\mathbf{\bar{x}}^{t})$. By using an argument similar to the derivation of (82), we have
@@
$\textstyle\leq\mathbb{E}[\|(\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t}))^{\top}\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|(\nabla\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1}))^{\top}\nabla F(\mathbf{\bar{x}}^{t-1})\|^{2}]$
$\textstyle\quad+2\mathbb{E}[\|(\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1})-\nabla\boldsymbol{g}(\boldsymbol{x}^{t}))^{\top}\nabla F(\mathbf{\bar{x}}^{t})\|]$
$\textstyle\quad+2\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\mathbf{\bar{x}}^{t-1})\|^{2}\|\nabla\boldsymbol{g}(\boldsymbol{x}^{t-1})\|^{2}]$
@@
$\textstyle\quad+2L_{g}^{2}L_{F}^{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}\|^{2}].$ (142)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*})\|^{2}]$
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$ (145)
$\textstyle\leq\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})-\nabla F(\boldsymbol{x}^{*})\|^{2}]$
@@
$\textstyle\mathbb{E}[F(\mathbf{\bar{x}}^{t+1})-F(\mathbf{\bar{x}}^{t})]\leq\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\mathbf{\bar{x}}^{t+1}-\mathbf{\bar{x}}^{t}\rangle]$ (146)
$\textstyle\quad+\frac{L_{F}^{2}}{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\mathbf{\bar{x}}^{t}\|^{2}].$
The first term on the right hand side of (146) satisfies
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\mathbf{\bar{x}}^{t+1}-\mathbf{\bar{x}}^{t}\rangle]=-\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\frac{\lambda_{x}^{t}}{m}\boldsymbol{u}^{t}\rangle]$
$\textstyle\leq\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]-\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\nabla F(\mathbf{\bar{x}}^{t})\rangle]$
$\textstyle\quad-\frac{\lambda_{x}^{t}}{m}\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\nabla F(\mathbf{\bar{x}}^{t})\rangle].$ (147)
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\boldsymbol{\breve{u}}^{t}-\nabla F(\mathbf{\bar{x}}^{t})\rangle]$ (148)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]+\frac{1}{\kappa}\mathbb{E}[\|\boldsymbol{\breve{u}}^{t}-\nabla F(\mathbf{\bar{x}}^{t})\|^{2}].$
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\breve{u}}^{t}-\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]$
$\textstyle\leq(2\bar{L}_{f}^{2}(1+L_{g}^{2})(1+2L_{g}^{2})+4\bar{L}_{g}^{2}L_{f}^{2})\mathbb{E}[\|\boldsymbol{x}^{t}-\mathbf{\bar{x}}^{t}\|^{2}]$
$\textstyle\leq\frac{c_{9}}{(t+1)^{2\varsigma_{x}}},$ (149)
where $c_{9}$ is given by $c_{9}=\bar{C}_{x}(2\bar{L}_{f}^{2}(1+L_{g}^{2})(1+2L_{g}^{2})+4\bar{L}_{g}^{2}L_{f}^{2})$ with $\bar{C}_{x}$ given in (44). Here, we have used $\mathbb{E}[\|\boldsymbol{x}^{t}-\mathbf{\bar{x}}^{t}\|^{2}]\leq\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\leq\frac{\bar{C}_{x}}{(t+1)^{2\varsigma_{x}}}$ in the last inequality.
@@
$\textstyle\mathbb{E}[\langle\nabla F(\mathbf{\bar{x}}^{t}),\mathbf{\bar{x}}^{t+1}-\mathbf{\bar{x}}^{t}\rangle]\leq-\frac{\lambda_{x}^{t}}{2m}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]$ (150)
$\textstyle\quad+\frac{b_{5}\lambda_{x}^{t}}{m\kappa(t+1)^{\beta_{1}}}+\frac{b_{6}^{\prime}\lambda_{x}^{t}}{m(t+1)^{\beta_{3}}}+\frac{c_{9}}{\kappa(t+1)^{2\varsigma_{x}}}.$
@@
$\mathbb{E}[F(\mathbf{\bar{x}}^{t+1})-F(\mathbf{\bar{x}}^{t})]\leq-\frac{\lambda_{x}^{t}}{2m}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]+\Phi_{t},$ (151)